OpenStudy (anonymous):

A cylinder has a radius of 14 m and a height of 6 m. What is the exact volume of the exact volume of the cylinder?

OpenStudy (anonymous):

A.1176pie m^3 b.504pie m^3 c.168pie m^3 d.84pie m^3

OpenStudy (anonymous):

@confluxepic

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

ill medal

OpenStudy (anonymous):

can you help @hockeychick23

OpenStudy (hockeychick23):

v=πr^2h

OpenStudy (anonymous):

i think its b.

OpenStudy (hockeychick23):

14^2=196*6= 1176π

OpenStudy (anonymous):

oooh so a. ty

OpenStudy (hockeychick23):

no problem :)

OpenStudy (anonymous):

can u help me with a few more ?@hockeychick23

OpenStudy (hockeychick23):

yea sure

OpenStudy (anonymous):

What is the exact volume of the cylinder? http://static.k12.com/calms_media/media/1422500_1423000/1422966/1/3bbab29ee2f972778fe6b61dc6a19e1b5da56204/MS_PA_131002_171325.jpg a.780pie in^3 b.390pie in^3 c5070pie in^3 d.11,700pie in^3

OpenStudy (anonymous):

@hockeychick23 i think its c

OpenStudy (hockeychick23):

ok so again v=πr^2h So 13^2= 169 and h=30 169*30= 5070 yes its C

OpenStudy (anonymous):

yay ty

OpenStudy (hockeychick23):

sure

OpenStudy (anonymous):

can u help with 2 more ? thats all i need ?

OpenStudy (hockeychick23):

k

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

What is the approximate volume of a cylinder with a height of 2 ft and a radius of 6 ft? Use 3.14 to approximate pi. Express your answer in hundredths. _______ft3

OpenStudy (hockeychick23):

v=πr^2h v= 6^2*2*3.14 v= 36*2*3.14 v=226.08 ft^3

OpenStudy (anonymous):

The volume of a cylinder, found using 3.14 to approximate pi, is 30,395.2 cm3. The radius of the cylinder is 11 cm. What is the height of the cylinder? Express your answer as a whole number. _________cm

OpenStudy (anonymous):

this is the last one @hockeychick23

OpenStudy (hockeychick23):

ok so v=πr^2h is still your formula so lets fill in the numbers: The volume of a cylinder, found using 3.14 to approximate pi, is 30,395.2 cm3. The radius of the cylinder is 11 cm. 30,395.2= 3.14*11^2*h 30,395.2= 3.14*121*h 30,395.2= 379.94*h Now divide both sides by 379.94 30,395.2/379.94= 80

OpenStudy (anonymous):

ooh ok so 80?

OpenStudy (hockeychick23):

Yes

OpenStudy (anonymous):

yay ty i got 100%

OpenStudy (hockeychick23):

awesome!(:

OpenStudy (anonymous):

that helped me alot xD ty

OpenStudy (hockeychick23):

anytime :)